Skip to end of metadata
Go to start of metadata

AHATS (sometime)daily status 8/3/08

Staff: Militzer, Oncley, Nguyen, Tudor&Verstraete with ISS.
John replaced Steve yesterday.  Lou arrived the day before and
Laura leaves tomorrow.

Temps: 99F/57F yesterday

Activity highlights:
- Changed sensor to monitor Pref to free up the fast-response sensor to
make a completely independent measurement.  All measurements appear to
agree...
- I noticed dust devils in the pressure signal yesterday -- these
probably are the first ever with a tight array of pressure measurements!
- We <think> we've completed documentation of array 3.

For more details, see the ahats logbook at https://wiki.ucar.edu/display/ahatslogbook

Good wind direction:
array#3 (jul 29 12:30-present) now has 40 hours (8 unstable, 32 stable).
We're planning on switching to array #4 on Tom's schedule on 6-7 Aug.

Local data storage: (we downloaded everything on 30 July)
------------------
upwind:/dev/sda1 57685532 7961492 49724040 14% /var/tmp/usbdisk
downwind1:/dev/sda1 57685532 8849776 48835756 15% /var/tmp/usbdisk
downwind2:/dev/sda1 57685532 10443500 47242032 18% /var/tmp/usbdisk
profile:/dev/sda1 57685532 5801844 51883688 10% /var/tmp/usbdisk
pressure1:/dev/sda1 3897204 1800328 2096876 46% /var/tmp/usbdisk
pressure2:/dev/sda1 3897204 1857176 2040028 48% /var/tmp/usbdisk

aster:/dev/sdb1 721075720 368028044 353047676 52% /media/isff2
isff: /dev/sdc1 1922890480 544537872 1280675404  30% /media/isff15

Pressure:
--------
(+/- = ~1 std deviation among variables at the same height)

11.b now measuring p with the 202BG connected between the
Bedard port and CHATS reference.  P.0m [poor name] now is the
portable Paroscientific standard with reasonable resolution, but
low sampling rate.

p: ok, +/- 0.03 mb
p'p': ok, +/- 0.0005 mb^2
w'p': ok, +/- 0.001 m/s mb
t: ok, 2 degC
Pref: ok, glitch yesterday ~noon when opened to install new sensor

P: ok

Profile: [,select.p]
--------
glitches on 8m when we replaced it today (0900-0930)

T: ok
RH: sensor that is now at 5.8m is strange.  Since sonics are now at this
level, we may work to fix it.

diag: ok
samples.sonic: ok
spd: ok?, 8m still lower than 7m sometimes.  (Even after the swap today!)
Maybe the WiFi antenna is bothering it...
dir: ok
w: ok
tc: ok, 8m low by ~0.3 degC? (related to the speeds being low as well?)
w'w': ok
u*: ok, some imaginary values with light winds
sigma_w/u*:
w'tc': ok
h2o: normal, 2 g/m^3 offset from dat("Q"), except in afternoon.
w'h2o': ok, 0.03 m/s g/m^3 at midday
co2: ok, 15-20 mmol/m^3 (some nights have large values, some don't [last
night]).  Presumably related to shallow nocturnal boundary layers,
but I haven't gone through the sodar data yet.
w'co2': ok, -0.01 m/s mmo/m^3 at midday

Upwind (hts=4.8): [,select.u]
----------------

diag: ok
samples.sonic: ok
spd: ok, +/- 20 cm/s
dir: ok, why isn't this better now that angles are shot??
w: ok, +/- 5 cm/s
tc:  ok, +/- 0.2 deg, biases somewhat larger
w'w': ok, +/- 0.005 m^2/s^2  (30 min avg for second moments)
u*: ok,  +/- 2 cm/s, some imaginary
sigma_w/u*:
w'tc': ok, +/- 0.01 m/s degC
tc'tc': ok, +/- 0.05 degC^2

Downwind Lower (hts=4.8): [,select.b]
------------------------

diag: ok
samples.sonic: ok

spd: ok, +/- 20 cm/s
dir: ok, why isn't this better now that angles are shot??
w: ok, +/- 5 cm/s, leveling helped!
tc:  ok, +/- 0.2 deg, biases up to +/-0.8
w'w': ok, +/- 0.005 m^2/s^2  (30 min avg for second moments)
u*: ok,  +/- 2 cm/s, some imaginary
sigma_w/u*:
w'tc': ok, +/- 0.01 m/s degC
tc'tc': ok, +/- 0.05 degC^2

Downwind Upper (hts=5.8): [,select.t]
------------------------

diag: ok
samples.sonic: ok
spd: ok, +/- 10 cm/s, 3t outlier
dir: ok, +/- 2 deg, why isn't this better now that angles are shot??
11t outlier
w: ok, +/- 5 cm/s, leveling helped!
tc:  ok, +/- 0.2 degC, offsets up to 0.6 degC
w'w': ok, +/- 0.01 m^2/s^2  (30 min avg for second moments)
u*: ok,  +/- 2 cm/s, som imaginary
sigma_w/u*:
w'tc': ok, +/- 0.01 m/s degC
tc'tc': ok, +/- 0.05 degC^2

----------------------------------------------------------------